# Multimodal Clinical Trial Outcome Prediction with Large Language Models

### v3

@@ 3. Experiments
3.2. Experimental Setup
Baselines. We compare LIFTED with both machine learning methods, including Logistic regression (LR) (Siah et al., 2021; Lo et al., 2019), Random Forest (RF) (Lo et al., 2019; Siah et al., 2021), XGBoost (Rajpurkar et al., 2020; Siah et al., 2021), Adaptive boosting (AdaBoost) (Fan et al., 2020), k Nearest Neighbor (kNN) + RF (Lo et al., 2019) and deep learning models, such as Feedforward Neural Network (FFNN) (Tranchevent et al., 2019), DeepEnroll (Zhang et al., 2020), COMPOSE (Gao et al., 2020), HINT (Fu et al., 2022; Chen et al., 2024), SPOT (Wang et al., 2023b). Among these, HINT and SPOT are specifically designed for clinical trial outcome prediction. HINT encodes and integrates different modalities with modality-specific encoders, including a drug molecule encoder utilizing MPNN algorithm, a disease ontology encoder based on GRAM, a trial eligibility criteria encoder leveraging BERT, and also, a drug molecule pharmacokinetic encoder, surplus a graph neural network to capture feature interactions. SPOT identifies trial topics firstly to group the diverse trial data from multiple sources into relevant trial topics, and then produces and organizes trial embeddings according to topic and timestamp to construct organized clinical trial sequences. Finally, Each trial sequence is treated as a separate task, and a meta-learning approach is employed to adapt to new tasks with minimal modifications. More details of those baselines are presented in Appendix B.
Evaluation Metrics. Following Fu et al. (2022) and Wang et al. (2023b), we use F1 score, PR-AUC, and ROC-AUC to measure the performance of all methods. For all these three metrics, higher scores indicate better performance.
Hyperparameter Settings. We follow the settings of most hyperparameters in HINT (Fu et al., 2022). The models are trained for a total of 5 epochs using a mini-batch size of 32 on one NVIDIA 4090 GPUs, which will take up to 2 hours. We employ the AdamW optimizer with a learning rate of $3\times 10^{-4}$, $\beta$ values of $(0.9,0.99)$, and a weight decay of $1\times 10^{-2}$ with a CosineAnnealing learning rate scheduler.
Table 3. The clinical trial outcome prediction performance (%) of LIFTED and variants without certain key component. The best results are bold. LIFTED outperforms all variants, showcasing the effectiveness of our proposed components.
@@ 3.2. Experimental Setup
Table 4. Performance analysis of multimodal data integration. The best results and second best results are bold and underlined, respectively.
@@
3.3. Overall Performance
We conduct experiments to evaluate the performance of LIFTED on phase I, II and III data respectively compared to our baselines. The trial outcome prediction results of all models are reported in Table 2. We first observed that the deep learning based methods, especially the methods designed for clinical trial outcome prediction including HINT and SPOT, outperforms the machine learning based methods with a significant performance gap, showcasing the powerful ability to extract critical information from different modalities in various formats of the deep learning encoders, especially those encoders specifically designed to extract representation hidden in the clinical trial records. This observation is not surprising, since the critical information of different modalities is represented in different ways, which is hard to extract for those traditional machine learning methods or those deep learning encoders that are not designed for clinical trial outcome prediction. Nevertheless, LIFTED consistently outperforms all other methods over all three phases, verifying its effectiveness in unifying different modalities and dynamically integrating them within the MoE.
3.4. Ablation Study
In this section, we perform comprehensive ablation studies to demonstrate the effectiveness of our key components, including the representation augmentation, the auxiliary loss and the modalities used to generate weights in the Mixture-of-Experts (MoE) framework. The ablation models are described as:
  • LIFTED-aug: In LIFTED-aug, the representation augmentation component and the consistency loss are removed. Representations from different modalities are directly fed into the multimodal data integration component without the constraint of robustness to the noise in the data.
  • LIFTED-aux: In LIFTED-aux, we remove the auxiliary loss component. Representations from different modalities are no longer required to make consistent predictions with the final representation integrated by the MoE framework.
3.5. Analysis of Multimodal Data Integration
We further analyze how multimodal data integration contributes to clinical outcome prediction. Here, we compare the performance of models using data from only one modality with LIFTED that integrates all those modalities. We report the results in Table 4. The results indicate that LIFTED outperforms all unimodal models in terms of all metrics, demonstrating the effectiveness of multimodal integration. The reason why the F1 score of LIFTED is lower than that of the drug description unimodal models on phase I and III is that the drug description unimodal models, tend to produce all-positive or all-negative predictions, which results in unexpectedly high F1 score with near zero standard deviation due to the high success rate on the dataset. In addition, the results also demonstrate that the drug description modality is the least important modality, while the criteria modality is the most important modality. This is within expectation since the quality of recruited patients plays a crucial role in trial success (Jin et al., 2017).
3.6. Analysis of Sparse Mixture-of-Experts
In addition, we delve into an analysis of the Sparse MoE model to understand the performance enhancements obtained by the sparse MoE model. Here, we select a knee osteoarthritis patient case. For each modality, the SMoE framework selects top-$3$ experts from a pool of $16$ experts with the highest weights. The weights of these selected SMoE experts are visualized in Figure 4. As expected, certain experts, such as 6 and 7, are consistently chosen across multiple modalities, indicating their pivotal role in extracting similar information patterns among different modalities. Furthermore, other experts demonstrate a more focused expertise, concentrating on one or two modalities. This demonstrates the effectiveness of the SMoE framework in both extracting similar information patterns across different modalities and capturing specialized information patterns within a single modality.
Figure 4. The SMoE experts’ importance weights of our model predicting the knee osteoarthritis patient. Experts 6 and 7 play a crucial role in extracting common information patterns across modalities, while other experts specialize in a single specific modality.
3.7. Case Study
In addition, we conduct a case study to analyze the contribution of each modality in clinical trial outcome prediction. Specifically, we analyze the result of a type 2 diabetes mellitus patient, who was inadequately controlled with metformin at the maximal effective and tolerated dose of metformin for at least 12 weeks. Since type 2 diabetes mellitus is hard to cure, the model should pay attention to the name of the disease and predict the trial as failed, which is consistent with the behavior of our model. The modality importance weights are shown in Figure 5. As we expected, the attention weights of the disease modality are much higher than other modalities, which demonstrates that our model pays attention to the disease modality and predicts the trial correctly.
Figure 5. The modality importance weights of our model predicting the type 2 diabetes mellitus patient. LIFTED pay more attention to the disease modality as expected, since type 2 diabetes mellitus is hard to cure.
4. Related Works
Clinical Trials Outcome Prediction. Machine learning algorithms have been proven efficient on diverse tabular data prediction tasks, including the clinical trial outcome prediction task, resulting in profound performances (Huang et al., 2020b; Arik and Pfister, 2021; Chen et al., 2022c; Klambauer et al., 2017; Badirli et al., 2020; Chen et al., 2023; Hazimeh et al., 2020; Chen et al., 2022b; Popov et al., 2019; Yan et al., 2023, 2024; Grinsztajn et al., 2022; Luo et al., 2023). For instance, Gayvert et al. (2016) employed random forests to integrate drug chemical structures and properties to predict drug toxicity; Qi and Tang (2019) utilized recurrent neural networks (RNNs) to predict pharmacokinetic outcomes at phase III, leveraging data acquired from phase II trials. Similarly, Lo et al. (2019) employed statistical machine learning models to predict drug approval. Recently, Fu et al. (2022, 2023) proposed a hierarchical interaction network employing different encoders to fuse multiple modal data and capture their correlations for trial outcome predictions; based on Fu et al. (2022), Chen et al. (2024) quantifies uncertainty in the prediction; Wang et al. (2023b) clustered multi-sourced trial data into different topics, organizing trial embeddings for prediction. Wang et al. (2023a) converted clinical trial data into a format compatible description for prediction. However, converting all modalities into a single description poses significant challenges. This approach makes it difficult for the model to distinguish the unique information of each modality and necessitates external data to aid in differentiating these modalities. In contrast, LIFTED extracts representations for each modality separately and dynamically integrates them, providing a more effective way to preserve distinct characteristics of each modality.
Mixture-of-Experts. Mixture-of-Experts (MoE) is a special type of neural network whose parameters are partitioned into a series of sub-modules, called experts, functioning in a conditional computation fashion (Jacobs et al., 1991; Jordan and Jacobs, 1993). Since traditional dense MoE models (Eigen et al., 2013) utilize all experts for each input, they are computationally expensive. Recently, Shazeer et al. (2017) simplified the MoE layer by selecting a sparse combination of the experts, instead of all experts, to process input data, significantly reducing the computational cost and improving the training stability. Subsequently, Fedus et al. (2022) further reduced the routing computation cost of the MoE layer by routing one sample to only a single expert instead of $K$ experts, enabling the scaling of language models to enormous sizes, such as trillions of parameters, without sacrificing performance. To encourage specialization and decrease redundancy among experts (Chen et al., 2022a), Dai et al. (2022) pre-defined the expert assignment for different input categories, and Hazimeh et al. (2021) advocated multiple, diverse router policies, facilitating the intriguing goals of SMoE is to divide and conquer the learning task by solving each piece of the task with adaptively selected experts (Aoki et al., 2022; Hazimeh et al., 2021; Ma et al., 2018; Mittal et al., 2022). In addition, different neural network structures (Dauphin et al., 2017; Vaswani et al., 2017) have been proposed and achieved surprising successes in various applications (Shahbaba and Neal, 2009; Lepikhin et al., 2020; Zhou et al., 2022; Dauphin et al., 2017; Riquelme et al., 2021; Eigen et al., 2013; Ahmed et al., 2016; Gross et al., 2017; Zhang et al., 2023; Yao et al., 2021; Wu et al., 2023). To identify similar information patterns between different modalities and extract them with the same expert model, LIFTED follows the original design of Sparse Mixture-of-Experts (Shazeer et al., 2017), routing inputs to a subset of experts instead of just one expert, dynamically selecting the experts instead of using a pre-defined assignment.
5. Conclusion
This paper presents LIFTED, an approach that unifies multimodal data using natural language descriptions and integrates this information within a Mixture-of-Experts (MoE) framework for clinical trial outcome prediction. We employ noise-resilient encoders to extract representations from each modality, utilize a Sparse Mixture-of-Experts framework to further dig the similar information patterns in different modalities and incorporate a dynamic representation integration approach with an MoE framework. Additionally, we introduce an auxiliary loss to improve the quality of modal-specific representations. Empirically, our LIFTED method demonstrates superior performance compared to existing approaches across all three phases of clinical trials.
Acknowledgements
We thank Google Cloud Research Credits program for supporting our computing needs.
Appendix A Prompt
The whole prompt, including the system message, is demonstrated in Table 5, and some examples are demonstrated in Table 6.
Table 5. Prompting.
System Message
You are a helpful assistant.
Prompting
Response
\$summarization_of_the_sample
Table 6. Examples of Prompting.
Linearization Summarization
[…]
Appendix B Baselines
Many methods have been selected as baselines in our experiments, including both statistical machine learning and deep learning models. We use the same setups in Fu et al. (2022) and Wang et al. (2023b) for most of them.
  • Logistic regression (LR) (Lo et al., 2019; Siah et al., 2021): logistic regression with the default hyperparameters implemented by scikit-learn (Pedregosa et al., 2011).
  • Random Forest (RF) (Lo et al., 2019; Siah et al., 2021): similar to logistic regression, the random forest is also implemented by scikit-learn with the default hyperparameters (Pedregosa et al., 2011).
  • XGBoost (Rajpurkar et al., 2020; Siah et al., 2021): An implementation of gradient-boosted decision trees optimized for speed and performance.
  • Adaptive boosting (AdaBoost) (Fan et al., 2020): an adaptive boosting-based decision tree method implemented by scikit-learn (Pedregosa et al., 2011).
  • k Nearest Neighbor (kNN) + RF (Lo et al., 2019): a combined model using kNN to imputate missing data and predicting by random forests.
  • Feedforward Neural Network (FFNN) (Tranchevent et al., 2019): a feedforward neural network that uses the same feature as HINT (Fu et al., 2022). The FFNN contains three fully-connected layers with hidden dimensions of 500 and 100, as well as a rectified linear unit (ReLU) activation layer to provide nonlinearity.
  • DeepEnroll (Zhang et al., 2020): initially intended for patient-trial matching, DeepEnroll employs three key components: (1) a pre-trained BERT model (Devlin et al., 2019) to encode eligibility criteria into sentence embeddings, (2) a hierarchical embedding model to handle disease information, and (3) an alignment model to capture interactions between eligibility criteria and diseases. In our experiments, to adapt DeepEnroll for predicting trial outcomes, its functionality is extended by concatenating the molecule embeddings (hm) computed by the MPNN algorithm (Huang et al., 2020a) over molecule graphs with the output of the alignment model.
  • COMPOSE (Gao et al., 2020): similar to DeepEnroll, COMPOSE was originally proposed for patient-trial matching. A convolutional neural network and a memory network are employed to encode eligibility criteria and diseases, respectively. Similarly, a molecule embedding from MPNN is also concatenated with its embedding for trial outcome prediction.
  • HINT (Fu et al., 2022): several key components are integrated with HINT, including a drug molecule encoder utilizing MPNN algorithm, a disease ontology encoder based on GRAM, a trial eligibility criteria encoder leveraging BERT, and also, a drug molecule pharmacokinetic encoder, surplus a graph neural network to capture feature interactions. After the interacted features are encoded, they are fed into a prediction model for accurate outcome predictions.
  • SPOT (Wang et al., 2023b): SPOT contains several steps. Firstly, trial topics are identified to group the diverse trial data from multiple sources into relevant trial topics. Subsequently, trial embeddings are produced and organized according to topic and timestamp to construct organized clinical trial sequences. Finally, Each trial sequence is treated as a separate task, and a meta-learning approach is employed to adapt to new tasks with minimal modifications.

### v4

@@ Appendix E Detailed Results
CTOD
@@
LR $55.6\pm 2.3$ $58.5\pm 2.1$ $51.5\pm 2.2$ $56.0\pm 1.4$ $63.6\pm 1.1$ $56.7\pm 1.0$ $75.0\pm 2.2$ $73.8\pm 1.6$ $51.9\pm 2.0$
RF $58.3\pm 2.9$ $70.0\pm 1.6$ $54.0\pm 2.0$ $60.8\pm 1.3$ $70.9\pm 0.8$ $56.0\pm 1.3$ $75.1\pm 1.6$ $85.1\pm 0.7$ $50.7\pm 2.2$
XGBoost $57.4\pm 3.5$ $68.0\pm 1.9$ $50.5\pm 2.3$ $61.2\pm 1.5$ $70.6\pm 1.2$ $61.3\pm 1.3$ $78.7\pm 1.8$ $84.9\pm 0.7$ $57.1\pm 2.0$
AdaBoost $58.8\pm 3.2$ $66.7\pm 2.4$ $53.5\pm 2.3$ $60.2\pm 1.8$ $68.9\pm 0.8$ $55.2\pm 1.3$ $79.5\pm 1.3$ $84.6\pm 0.8$ $56.9\pm 1.9$
kNN+RF $58.4\pm 3.9$ $70.3\pm 2.0$ $58.7\pm 2.1$ $61.2\pm 1.7$ $70.9\pm 0.9$ $56.6\pm 1.5$ $76.8\pm 2.2$ $85.2\pm 1.1$ $52.5\pm 1.9$
FFNN $55.2\pm 2.2$ $58.4\pm 1.4$ $48.7\pm 1.7$ $57.2\pm 1.5$ $66.0\pm 1.3$ $53.0\pm 1.2$ $76.1\pm 1.4$ $79.6\pm 1.1$ $52.6\pm 1.5$
MMF (early fusion) $61.2\pm 2.8$ $64.2\pm 2.1$ $56.2\pm 2.8$ $64.0\pm 1.2$ $70.1\pm 0.9$ $59.0\pm 1.1$ $83.3\pm 1.4$ $84.4\pm 1.0$ $64.6\pm 1.9$
MMF (late fusion) $63.0\pm 3.1$ $70.5\pm 1.8$ $57.4\pm 2.4$ $65.7\pm 2.1$ $71.4\pm 0.8$ $60.3\pm 2.6$ $83.5\pm 1.7$ $85.2\pm 0.8$ $65.8\pm 1.6$
HINT $63.4\pm 2.9$ $71.0\pm 1.2$ $57.6\pm 2.4$ $64.6\pm 2.4$ $71.2\pm 1.3$ $60.6\pm 2.0$ $81.8\pm 1.9$ $85.7\pm 0.7$ $60.8\pm 2.0$
SPOT $66.8\pm 1.9$ $70.0\pm 0.9$ $62.3\pm 1.0$ $64.5\pm 1.7$ $71.8\pm 0.7$ $58.8\pm 1.4$ $83.2\pm 2.3$ $77.6\pm 1.0$ $67.5\pm 1.7$
LIFTED (ours) 69.7 $\pm$ 3.0 71.8 $\pm$ 1.4 63.4 $\pm$ 2.3 67.7 $\pm$ 2.0 72.0 $\pm$ 1.4 63.0 $\pm$ 1.2 86.7 $\pm$ 1.2 85.9 $\pm$ 1.0 69.5 $\pm$ 1.8
Table E.5: The clinical trial outcome prediction performance (%) of LIFTED and variants without certain key component. The best results are bold. LIFTED outperforms all variants, showcasing the effectiveness of our proposed components.
HINT
Phase I Trials Phase II Trials Phase III Trials
Method PR F1 ROC PR F1 ROC PR F1 ROC
@@
CTOD
Phase I Trials Phase II Trials Phase III Trials
Method PR F1 ROC PR F1 ROC PR F1 ROC
LIFTED-aug $65.1\pm 3.7$ $70.6\pm 1.7$ $61.9\pm 2.4$ $67.1\pm 1.8$ $70.5\pm 1.1$ $62.2\pm 1.4$ $85.2\pm 1.4$ $82.4\pm 1.1$ $67.8\pm 1.6$
LIFTED-aux $64.4\pm 3.3$ $71.1\pm 1.8$ $60.0\pm 2.8$ $60.0\pm 1.4$ $71.0\pm 0.9$ $54.9\pm 1.0$ $83.8\pm 1.8$ $85.5\pm 0.9$ $64.6\pm 2.4$
LIFTED-LLM $65.1\pm 2.7$ $71.4\pm 1.6$ $58.5\pm 3.2$ $66.3\pm 1.9$ $71.1\pm 1.1$ $61.2\pm 1.5$ $83.9\pm 1.3$ $85.4\pm 1.0$ $65.3\pm 2.0$
LIFTED-gating 67.1 $\pm$ 2.4 71.4 $\pm$ 1.8 60.8 $\pm$ 2.5 65.1 $\pm$ 1.4 70.4 $\pm$ 1.3 61.9 $\pm$ 1.5 85.0 $\pm$ 1.5 85.8 $\pm$ 0.7 68.8 $\pm$ 1.8
LIFTED (ours) 69.7 $\pm$ 3.0 71.8 $\pm$ 1.4 63.4 $\pm$ 2.3 67.7 $\pm$ 2.0 72.0 $\pm$ 1.4 63.0 $\pm$ 1.2 86.7 $\pm$ 1.2 85.9 $\pm$ 1.0 69.5 $\pm$ 1.8
Table E.6: Performance analysis of multimodal data integration. The best results and second best results are bold and underlined, respectively.
HINT
@@
CTOD
Phase I Trials Phase II Trials Phase III Trials
PR-AUC F1 ROC-AUC PR-AUC F1 ROC-AUC PR-AUC F1 ROC-AUC
Summarization $61.6\pm 3.2$ $70.9\pm 2.2$ $56.9\pm 2.6$ $65.4\pm 1.3$ $71.5\pm 0.9$ $61.1\pm 1.3$ $84.1\pm 1.3$ $84.8\pm 0.9$ $64.9\pm 1.9$
Drugs $60.8\pm 2.8$ $70.6\pm 1.4$ $57.6\pm 2.4$ $58.0\pm 1.8$ $71.3\pm 1.1$ $53.8\pm 1.2$ $77.3\pm 1.7$ $85.2\pm 1.0$ $54.5\pm 1.9$
Disease $65.2\pm 2.0$ $70.2\pm 1.9$ $61.3\pm 1.7$ $68.3\pm 1.5$ $71.1\pm 1.2$ $62.6\pm 1.0$ $85.3\pm 1.4$ $85.5\pm 0.7$ $68.4\pm 2.0$
Description $56.6\pm 2.6$ $70.7\pm 1.5$ $52.3\pm 2.7$ $56.0\pm 1.8$ $71.3\pm 0.8$ $52.4\pm 1.7$ $76.5\pm 1.2$ 85.7 $\pm$ 0.7 $52.9\pm 1.9$
SMILES $61.6\pm 2.8$ $70.6\pm 1.7$ $56.0\pm 2.4$ $58.0\pm 1.8$ $71.5\pm 1.2$ $53.1\pm 1.2$ $75.1\pm 1.6$ $85.6\pm 0.7$ $50.8\pm 2.4$
Criteria $60.5\pm 3.0$ $71.2\pm 1.7$ $55.8\pm 3.1$ $59.9\pm 2.5$ 71.2$\pm$ 1.2 $53.9\pm 2.0$ $75.7\pm 2.0$ $84.6\pm 0.7$ $52.5\pm 2.6$
LIFTED (ours) 69.7 $\pm$ 3.0 71.8 $\pm$ 1.4 63.4 $\pm$ 2.3 67.7 $\pm$ 2.0 72.0 $\pm$ 1.4 63.0 $\pm$ 1.2 86.7 $\pm$ 1.2 85.9 $\pm$ 1.0 69.5 $\pm$ 1.8
The detailed experiments results with variants are presented in Table E.4, E.5, E.6